OpenStudy (anonymous):

Let f(x) = ln(e + 3x) Find the taylor series of the function f(x) centered at b=0. Could someone help please?

OpenStudy (dan815):

dannngg broo musscless

OpenStudy (dan815):

http://www.youtube.com/watch?v=cjPoEZ0I5wQ

OpenStudy (dan815):

everything u need to know about taylor and mclauren

OpenStudy (experimentx):

\[ \ln(e + 3x) = \ln(e (1 + 3x/e)) = \ln e + \ln (1 + 3x/e) = 1 + \sum_{n=1}^\infty (-1)^{n+1} (3x/e)^n\]